Determine the charge density due to each of the following electric flux densities: TI.
Answer #1

Maxwells first equations gives the relationship between charge density (P) and electric flux density (D) as follow: D. I = P
1. (b) i = up caso a9 + 24 sind ag + 2z² aq elme » P = 0. (in cylindrical Cordinate system) = Šl84p cas b) + Ø (29 Sind) + Q2

Sin 0 A also C/m2 XJ * l = 0.0 I (in spherical coordinate system) Put B ( 14 time (-a ono sing) smo 12616 & sino do et sine (